OpenStudy (anonymous):

A box with an open top has vertical sides, a square bottom, and a volume of 256 cubic meters. If the box has the least possible surface area, find its dimensions.

OpenStudy (aaronq):

 let h be the height and x the side lengths \(V=h*x^2=256~m^3\rightarrow h=\dfrac{256~m}{x^2}\) \(SA=x^2+4hx\) If we only count the outside faces. \(SA=2x^2+8hx\) If we count the inside faces as well as the outside faces. Well go with the second one, sub in h \(SA=2x^2+8(\dfrac{256}{x^2})x=2x^2+\dfrac{2048}{x}\) Now, we take the first derivative and let it equal to zero. \(SA'=4x-\dfrac{2048}{x^2}=0\) Solve for x and find h with the initial equation.